OpenStudy (anonymous):

I need help with an integral. It's part of an arc-length problem I'm working with but am stuck. I have the answer if that helps but am not sure how to get there. Integral from 0 to 2 of the square root of (t^2 + 16t +16) dt The answer in the back is (10root13 -16) - 24 ln((5+root13)/6). would really appreciate any advice. thanks.

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\sqrt{t ^{2}+16t+16}dt\]

OpenStudy (anonymous):

Answer:\[(10\sqrt{13}-16)-24\ln ((5+\sqrt{13})/6)\]

OpenStudy (anonymous):

Have you tried completing the square under the radical?

OpenStudy (anonymous):

so that would give me the square root of (t+8)^2 -48, right? then a trig substitution or something?

OpenStudy (anonymous):

Yes, just make sure you make the right sub.

OpenStudy (anonymous):

ok, cool. will give it a shot. thanks!